Kubin Companys relevant range of production is 11,000 to 14,000 units. When it produces and sells 12.500 units, its average costs per unit are as follows: Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense per Uni $ 7.20 $ 4.20 $ 1.70 $ 4.20 S3.70 $ 2.70 1.20 $ 0.70 Sales commissions Variable administrative expense Required: 1. If 11,000 units are produced and sold, what is the variable cost per unit produced and sold? 2. If 14,000 units are produced and sold, what is the variable cost per unit produced and sold? 3. If 11,000 units are produced and sold, what is the total amount of variable cost related to the units produced and sold? 4. If 14,000 units are produced and sold, what is the total amount of variable cost related to the units produced and sold? 5. If 11,000 units are produced, what is the average fixed manufacturing cost per unit produced? 6. If 14,000 units are produced, what is the average fixed manufacturing cost per unit produced? 7. If 11,000 units are produced, what is the total amount of fixed manufacturing overhead incurred to support this level of production? 8. If 14,000 units are produced, what is the total amount of fixed manufacturing overhead incurred to support this level of production? (Round per unit values to 2 decimal places.)
Required: 1. If 11,000 units are produced and sold, what is the variable cost per unit produced and sold? 2. If 14,000 units are produced and sold, what is the variable cost per unit produced and sold? 3. If 11,000 units are produced and sold, what is the total amount of variable cost related to the units produced and sold? 4. If 14,000 units are produced and sold, what is the total amount of variable cost related to the units produced and sold? 5. Ir 11,000 units are produced, what is the average fixed manufacturing cost per unit produced? 6. If 14,000 units are produced, what is the average fixed manufacturing cost per unit produced? 7 If 11,000 units are produced, what is the total amount of fixed manufacturing overhead incurred to support this level of production? 8. If 14,000 units are produced, what is the total amount of fixed manufacturing overhead incurred to support this level of production? (Round per unit values to 2 decimal places.) S 15.00 S 15.00 $165.000 $210 1 Variable cost per unit 2. Variable cost per unit 3. Total amount of variable cost 4. Total amount of variable cost 5. Average fixed manufacturing cost per unit 6. Average fixed 7 Total amount of fixed manufacturing overhead 8. Total amount of fixed manufacturing overhead cost per unit
Answer #1
1
Direct materials 7.20
Direct labor 4.20
Variable manufacturing overhead 1.70
Sales commissions 1.20
Variable administrative expense 0.70
Variable cost per unit produced and sold 15.00
2
Direct materials 7.20
Direct labor 4.20
Variable manufacturing overhead 1.70
Sales commissions 1.20
Variable administrative expense 0.70
Variable cost per unit produced and sold 15.00
3
Unit produced and sold 11000
X Variable cost per unit 15.00
Total amount of variable cost 165000
4
Unit produced and sold 14000
X Variable cost per unit 15.00
Total amount of variable cost 210000
5
Fixed manufacturing costs 52500 =12500*4.20
Divide by units 11000
Average fixed manufacturing cost per unit 4.77
6
Fixed manufacturing costs 52500 =12500*4.20
Divide by units 14000
Average fixed manufacturing cost per unit 3.75
7
Total amount of fixed manufacturing overhead 52500 =12500*4.20
8
Total amount of fixed manufacturing overhead 52500 =12500*4.20